# Benchmarking stabilized and self-stabilized $p$-virtual element methods with variable coefficients

### v1

@@ 4 Linear Elasticity problem
The local virtual element space, with order of accuracy $p\geq 1$ is defined for $p\leq m$, as [mengolini2019engineering]:
$\bm{\mathcal{V}}_{h}\left(E\right)=\left\{\right$. $\bm{u}_{h}\in\left[H^{1}\left(E\right)\cap C^{0}\left(E\right)\right]^{2}:\bm{L}\bm{\sigma}\left(\bm{u}_{h}\right)\big|_{E}\in\left[\mathcal{P}_{p-2}\left(E\right)\right]^{2},$ (44)
@@
where the operator $\bm{L}$ is the differential matrix defined as:
$\bm{L}=\begin{bmatrix},x&0&,y\\ 0&,y&,x\end{bmatrix}$. (45)
@@
$P_{0}^{E}\left(\bm{\psi}_{i},\bm{q}\right)=\frac{1}{n_{v}}\sum_{j=1}^{2n_{v}}\text{dof}_{j}\left(\bm{\psi}_{i}\right)\text{dof}_{j}\left(\bm{q}\right)$ (49)
@@
where $c_{i}$ are constants representing the planar rigid body motions. The stabilization forms listed in Section 3.1 are still valid options for the problem under consideration. In this case, $\bm{S}^{E}_{1}$, $\bm{S}^{E}_{2}$, $\bm{S}^{E}_{4}$ and $\bm{S}^{E}_{5}$ should be multiplied by the trace of the consistency matrix [mengolini2019engineering].
@@
$\bm{\varepsilon}^{\mathrm{\times}}=\begin{bmatrix},y&0\\ 0&-,x\\ -,x&,y\end{bmatrix}$. (51)
Lastly, the formulation $\mathrm{V6}$ uses the same space of $\mathrm{V4}$ with polynomial projection onto the space
@@ 5 Stokes problem
where $\nu$ is the viscosity and $L_{0}^{2}\left(\Omega\right)$ is the subspace of $L^{2}\left(\Omega\right)$ of null-mean valued functions. In this work, the divergence-free approximation detailed in [daveiga2017divergence, dassi2020bricks] is considered. Therefore, for $p\geq 2$ and $p\leq m$, the local velocity space is defined as:
@@ 6 Numerical Results: Benchmarking stabilized and self-stabilized VEM
It is well known that the stabilization term may alter the results for high-order approximations, giving sub-optimal convergence [daveiga2016basic]. As the $p$-VEM is the focus of the present investigation, this section aims at clarifying the influence of the stabilization parameter and stabilization techniques on the accuracy and stability of the method.
@@
For the polynomial basis $\mathcal{Q}_{p}\left(E\right)$, the Modified Gram-Schmidt (MGS) orthonormalization [mascotto2018ill] is used to obtain $L^{2}\left(E\right)$ orthonormal polynomials. Indeed, they are more suitable in a $p$-VEM context than standard monomials, which suffer from numerical instabilities.
@@ 6.1.2 Parametric study on the tolerance of the rank estimation
The tolerance is varied starting from the Matlab default value
$\mathrm{tol}_{0}=\text{max}\left(\text{size}\left(\bm{K}^{E}\right)\right)\times\epsilon\left(\left\lVert\bm{K}^{E}\right\rVert\right)$}
and progressively increased by powers of ten. As a reference, the self-stabilized formulation $\mathrm{V3}$ [berrone2025stabilization] is considered.
@@
The maximum $\ell_{E}$ in Figure 5(c) corresponds to the central element. Indeed, in general, the higher the number of edges, the higher the polynomial order to guarantee full rank [berrone2025stabilization].
@@ 6.2 Linear elasticity problem
The analysis is conducted on the quadrilateral and on the Bézier-edge meshes, with order $p=1,\dots,10$. The comparison between the stabilized and self-stabilized formulations is presented in terms of accuracy by setting $\tau=0.5$ – following the guidelines of Ref. [artioli2017arbitrary]) – and $\mathrm{tol}=\mathrm{tol}_{0}$.
@@ 7 Equations with variable coefficients
In virtual element setting, the issue of dealing with variable coefficients has been addressed in [daveiga2016virtual] for second-order elliptic problems: the authors discussed how the choice of polynomial projection ($\Pi^{\nabla}_{p}\psi_{i}$ or $\Pi^{0}_{p-1}\nabla\psi_{i}$) may lead to suboptimal results. Moreover, the stabilization term, which is only scaled by a measure of the coefficient, may further alter the results. This aspect is one of the motivations that justify the introduction of self-stabilized formulations.
A new strategy is proposed here to capture the influence of variable coefficients in both stabilized and self-stabilized VEM. The idea is to construct the polynomial projection in terms of the bilinear form, even if a generally non-polynomial coefficient is involved. This is done by enhancing the classical VEM space and considering high-order quadrature rules. The acronym $\mathrm{VC}$-VEM is employed when referring to this approach.
@@ 7.1 Second-order elliptic problem
Assume that the coefficient $\mathcal{A}\left(x,y\right)$ is uniformly positive definite and smooth within each element of the mesh. The discretization of the above problem in the standard virtual element space $\mathcal{V}_{h}$ of Eq. (8) yields two admissible definitions of the local discrete bilinear form depending on the choice of the polynomial projector, namely:
@@
In the case of the Laplace equation, both constructions are robust, whenever $\mathcal{A}$ is constant. This is not anymore true in the case of variable coefficients: the discrete problem built upon $\Pi^{\nabla}_{p}$ is subject to a strong deterioration in the order of convergence for $p\geq 3$, as demonstrated in [daveiga2016virtual]. Thus, the use of the $\Pi^{0}_{p-1}$ was recommended for high-order schemes.
As mentioned, a new projection operator is proposed here where $\mathcal{A}\left(x,y\right)$ is explicitly taken into account. To this aim, the following enhanced space is defined starting from $\mathcal{V}_{h}$:
$\mathcal{{V}}^{\mathcal{A}}_{h}\left(E\right)=\left\{\right.u_{h}\in\mathcal{V}_{h}\left(E\right):$ (68)
$\left(u_{h},\text{div}\left(\mathcal{A}\left(x,y\right)\nabla q\right)\right)_{E}\left.=\left(\Pi^{0}_{p}u_{h},\text{div}\left(\mathcal{A}\left(x,y\right)\nabla q\right)\right)_{E}\;\forall q\in\mathcal{P}_{p}\left(E\right)\right\}.$
Note that the expression above requires the partial derivatives of $\mathcal{A}\left(x,y\right)$ to be available.
Then, for the stabilized virtual element, the new coefficient dependent projection is defined as:
$\begin{cases}\begin{aligned} &a^{E}\left(\Pi^{\mathcal{A}}_{p}\psi_{i},q\right)=a^{E}\left(\psi_{i},q\right)&&\forall q\in\mathcal{P}_{p}\left(E\right),\\ &P_{0}^{E}\left(\Pi^{\mathcal{A}}_{p}\psi_{i},q\right)=P_{0}^{E}\left(\psi_{i},q\right)&&\forall q\in\mathcal{P}_{0}\left(E\right),\end{aligned}\end{cases}$ (69)
for all $i=1,...,N_{d}\big|_{E}$. The left-hand side, which involves polynomials and the coefficient $\mathcal{A}\left(x,y\right)$, can be approximated by standard high-order quadrature rules. On the other hand, the right-hand side is computed by applying integration by parts and the enhancement condition in the definition of the discrete space, hence:
$a^{E}\left(\psi_{i},q\right)=$ $-\int_{E}\Pi^{0}_{p}\psi_{i}\text{div}\left(\mathcal{A}\left(x,y\right)\nabla q\right)\;\,\mathrm{d}E$ (70)
@@
Once $\Pi^{\mathcal{A}}_{p}$ is computed, the discrete bilinear form is defined, as usual for VEM, by substituting the projection into the bilinear form and considering a suitable stabilization term (built on the standard $\Pi^{\nabla}_{p}$).
Concerning the self-stabilized formulations, the discrete space for $\mathrm{V1}$, $\mathrm{V3}$, $\mathrm{V4}$ and $\mathrm{V6}$ is obtained by enhancing the related spaces defined in Section 3.2 with the additional condition involving the coefficient $\mathcal{A}\left(x,y\right)$. This condition must be modified accordingly to the choice of the higher-order polynomial space. Moreover, in lieu of the $L^{2}$ projection, the following operator is considered:
@@
Note, $\mathrm{VC}$-VEM does not apply to formulations $\mathrm{V2}$ and $\mathrm{V5}$ since the enhancement through $\mathcal{A}\left(x,y\right)$ replaces the role of the additional degrees of freedom.
@@ 7.2 Linear Elasticity problem
The VC-VEM is now extended to the linear elasticity problem defined in Eq. (42). As presented in the previous example, the elasticity space in Eq. (44) is modified as follows:
$\bm{\mathcal{{V}}}^{\mathbb{C}}_{h}\left(E\right)=\left\{\bm{u}_{h}\in\bm{\mathcal{V}}_{h}\left(E\right):\right$. (72)
$\left.\left(\bm{u}_{h},\left[\bm{L}\mathbb{C}\left(x,y\right)\bm{\varepsilon}\left(\bm{q}\right)\right]\right)_{E}=\left(\Pi^{0}_{p}\bm{u}_{h},\left[\bm{L}\mathbb{C}\left(x,y\right)\bm{\varepsilon}\left(\bm{q}\right)\right]\right)_{E}\;\forall\bm{q}\in\left[\mathcal{P}_{p}\left(E\right)\right]^{2}\right\},$
so that the following coefficient dependent projection is computable:
$\begin{cases}\begin{aligned} &a^{E}\left(\bm{\Pi}^{\mathbb{C}}_{p}\bm{\psi}_{i},\bm{q}\right)=a^{E}\left(\bm{\psi}_{i},\bm{q}\right)&&\forall\bm{q}\in\left[\mathcal{P}_{p}\left(E\right)\right]^{2},\\ &P_{0}^{E}\left(\bm{\Pi}^{\mathbb{C}}_{p}\bm{\psi}_{i},\bm{q}\right)=P_{0}^{E}\left(\bm{\psi}_{i},\bm{q}\right)&&\forall\bm{q}\in RM\left(E\right),\end{aligned}\end{cases}$ (73)
for $i=1,...,N_{d}\big|_{E}$, where now:
$a^{E}\left(\bm{\psi}_{i},\bm{q}\right)=\int_{E}\bm{\varepsilon}\left(\bm{\psi}_{i}\right)^{T}\mathbb{C}\left(x,y\right)\bm{\varepsilon}\left(\bm{q}\right)\,\mathrm{d}E.$ (74)
@@
The projection operators $\Pi^{\mathcal{A}}_{p}$ and $\bm{\Pi}^{\mathbb{C}}_{p}$ must be recomputed whenever the coefficient changes. This also applies when a mesh of uniform elements is considered, as the coefficient varies spatially. In principle, the $\mathrm{VC}$-VEM may result in a higher computational cost. However it is mitigated by the fact that the $p$-version of VEM is generally defined on coarse meshes.
@@ 8.1.1 Example 1: polynomial coefficients
First, the quadrilateral mesh is considered. The results are shown in Figure 15. The accuracy of the solution, as seen from Figure 15(a), is the same for all the simulations up to the order $p=4$. However, when $p>4$ some curves exhibit a diverging response. It is worth noting the excellent accuracy displayed by all the methods based on the newly introduced approach to consider coefficient-variability, even for $p\geq 5$. In this scenario, also the self-stabilized $\mathrm{V3}$ and the stabilized VEM with $\Pi^{0}_{p-1}$ projection demonstrate excellent convergence. This behavior is attributed to the use of the $L^{2}$ projection, which tends to exhibits better accuracy in the presence of variable coefficients, as already discussed in [daveiga2016virtual]. On the other hand, the formulation $\mathrm{V6}$, which does not consider variable coefficients, behaves as the stabilized VEM with $\Pi^{\nabla}_{p}$. Regarding the stiffness matrix conditioning, reported in Figure 15(b), no differences can be noted between the cases with or without variable coefficient projection.

### v2

@@ 4 Linear Elasticity problem
The local virtual element space, with order of accuracy $p\geq 1$ is defined for $p\leq m$, as [30]:
$\bm{\mathcal{V}}_{h}\left(E\right)=\left\{\right$. $\bm{u}_{h}\in\left[H^{1}\left(E\right)\cap C^{0}\left(E\right)\right]^{2}:\bm{L}\left[\bm{\sigma}\left(\bm{u}_{h}\right)\right]\big|_{E}\in\left[\mathcal{P}_{p-2}\left(E\right)\right]^{2},$ (44)
@@
where the operator $\bm{L}\left[\cdot\right]$ is the differential matrix defined as:
$\bm{L}\left[\cdot\right]=\begin{bmatrix},x&0&,y\\ 0&,y&,x\end{bmatrix}$. (45)
@@
$P_{0}^{E}\left(\bm{\psi}_{i},\bm{q}\right)=\frac{1}{n_{v}}\sum_{j=1}^{2n_{v}}\text{dof}_{j}\left(\bm{\psi}_{i}\right)\text{dof}_{j}\left(\bm{q}\right),$ (49)
@@
where $c_{i}$ are constants representing the planar rigid body motions. The stabilization forms listed in Section 3.1 are still valid options for the problem under consideration. In this case, $\bm{S}^{E}_{1}$, $\bm{S}^{E}_{2}$, $\bm{S}^{E}_{4}$ and $\bm{S}^{E}_{5}$ should be multiplied by the trace of the consistency matrix [30].
@@
$\bm{\varepsilon}^{\mathrm{\times}}\left(\cdot\right)=\begin{bmatrix},y&0\\ 0&-,x\\ -,x&,y\end{bmatrix}$. (51)
Lastly, the formulation $\mathrm{V6}$ uses the same space of $\mathrm{V4}$ with polynomial projection onto the space:
@@ 5 Stokes problem
where $\nu$ is the viscosity and $L_{0}^{2}\left(\Omega\right)$ is the subspace of $L^{2}\left(\Omega\right)$ of null-mean valued functions. In this work, the divergence-free approximation detailed in [12, 26] is considered. Therefore, for $p\geq 2$ and $p\leq m$, the local velocity space is defined as:
@@ 6 Numerical Results: Benchmarking stabilized and self-stabilized VEM
It is well known that the stabilization term may alter the results for high-order approximations, giving sub-optimal convergence [9]. As the $p$-VEM is the focus of the present investigation, this section aims at clarifying the influence of the stabilization parameter and stabilization techniques on the accuracy and stability of the method.
@@
For the polynomial basis $\mathcal{Q}_{p}\left(E\right)$, the Modified Gram-Schmidt (MGS) orthonormalization [28] is used to obtain $L^{2}\left(E\right)$ orthonormal polynomials. Indeed, they are more suitable in a $p$-VEM context than standard monomials, which suffer from numerical instabilities.
@@ 6.1.2 Parametric study on the tolerance of the rank estimation
The tolerance is varied starting from the Matlab default value:
$\mathrm{tol}_{0}=\text{max}\left(\text{size}\left(\bm{K}^{E}\right)\right)\times\epsilon\left(\left\lVert\bm{K}^{E}\right\rVert\right)$},
and progressively increased by powers of ten. As a reference, the self-stabilized formulation $\mathrm{V3}$ [16] is considered.
@@
The maximum $\ell_{E}$ in Figure 5(c) corresponds to the central element. Indeed, in general, the higher the number of edges, the higher the polynomial order to guarantee full rank [16].
@@ 6.2 Linear elasticity problem
The analysis is conducted on the quadrilateral and on the Bézier-edge meshes, with order $p=1,\dots,10$. The comparison between the stabilized and self-stabilized formulations is presented in terms of accuracy by setting $\tau=0.5$ – following the guidelines of Ref. [3]) – and $\mathrm{tol}=\mathrm{tol}_{0}$.
@@ 7 Equations with variable coefficients
In virtual element setting, the issue of dealing with variable coefficients has been addressed in [7] for second-order elliptic problems: the authors discussed how the choice of the polynomial projection ($\Pi^{\nabla}_{p}\psi_{i}$ or $\Pi^{0}_{p-1}\nabla\psi_{i}$) may lead to suboptimal results. Moreover, the stabilization term, which is only scaled by a measure of the coefficient, may further alter the results. This aspect is one of the motivations that justify the introduction of self-stabilized formulations.
A new strategy is proposed here to capture the influence of variable coefficients in both stabilized and self-stabilized VEM. The idea is to construct a polynomial approximation of virtual functions that combines the standard $\Pi^{0}_{p}$ and the features of the coefficient. This new strategy is built upon well-established virtual element spaces. The acronym $\mathrm{VC}$-VEM is hereinafter employed to refer to this approach.
@@ 7.1 Second-order elliptic problem
Assume that the coefficient $\mathcal{A}\left(x,y\right)$ is uniformly positive definite and smooth within each element of the mesh.
Two different discretizations of the above problem were proposed and studied in the framework of the standard virtual element space $\mathcal{V}_{h}$ introduced in Eq. (8), depending on the choice of the polynomial projector, namely:
@@
In the case of the Laplace equation, both constructions are robust, whenever $\mathcal{A}$ is constant. This is not anymore true in the case of variable coefficients: the discrete problem built upon $\Pi^{\nabla}_{p}$ is subject to a strong deterioration in the order of convergence for $p\geq 3$, as demonstrated in [7]. Thus, the use of the $\Pi^{0}_{p-1}$ was recommended for high-order schemes.
As mentioned, a new projection of $\mathcal{V}_{h}\left(E\right)$ onto $\mathcal{P}_{p}\left(E\right)$ is proposed here, where $\mathcal{A}\left(x,y\right)$ is explicitly taken into account. The stabilized virtual element method is constructed by considering, for $i=1,...,N_{d}\big|_{E}$, the polynomial approximation $\Pi^{\mathcal{A}}_{p}\psi_{i}$ of $\psi_{i}$ defined as the solution of:
$\begin{cases}\begin{aligned} &a^{E}\left(\Pi^{\mathcal{A}}_{p}\psi_{i},q\right)=\mathfrak{a}_{\mathsf{vc}}^{E}\left(\psi_{i},q\right)&&\forall q\in\mathcal{P}_{p}\left(E\right),\\ &P_{0}^{E}\left(\Pi^{\mathcal{A}}_{p}\psi_{i},q\right)=P_{0}^{E}\left(\psi_{i},q\right)&&\forall q\in\mathcal{P}_{0}\left(E\right),\end{aligned}\end{cases}$ (68)
where $\mathfrak{a}_{\mathsf{vc}}^{E}$ is defined as:
$\mathfrak{a}_{\mathsf{vc}}^{E}\left(\psi_{i},q\right)\coloneqq$ $-\int_{E}\Pi^{0}_{p}\psi_{i}\,\text{div}\left(\mathcal{A}\left(x,y\right)\nabla q\right)\;\,\mathrm{d}E$ (69)
@@
The term $\mathfrak{a}_{\mathsf{vc}}^{E}\left(\psi_{i},q\right)$ is a reasonable approximation of $a^{E}\left(\psi_{i},q\right)$ obtained by applying integration by parts and replacing $\psi_{i}$ with $\Pi^{0}_{p}\psi_{i}$ in the non-computable volume integral. The expression above requires the partial derivatives of $\mathcal{A}\left(x,y\right)$ to be available.
All quantities in (68), which involve polynomials and the coefficient $\mathcal{A}\left(x,y\right)$, can be computed by considering standard high-order quadrature rules.
Once $\Pi^{\mathcal{A}}_{p}$ is computed, the discrete bilinear form is defined, as usual for VEM, by substituting the projection into the bilinear form and considering a suitable stabilization term (built on the standard $\Pi^{\nabla}_{p}$), that is:
$a_{h}^{E}\left(\psi_{i},\psi_{j}\right)=a^{E}\left(\Pi^{\mathcal{A}}_{p}\psi_{i},\Pi^{\mathcal{A}}_{p}\psi_{j}\right)+\tau S^{E}\left(\left(I-\Pi^{\nabla}_{p}\right)\psi_{i},\left(I-\Pi^{\nabla}_{p}\right)\psi_{j}\right).$ (70)
Since $\Pi^{\mathcal{A}}_{p}$ is not an orthogonal projection, the (non-computable) terms of kind $a^{E}\left(\Pi^{\mathcal{A}}_{p}\psi_{i},\left(I-\Pi^{\mathcal{A}}_{p}\right)\psi_{j}\right)$ do not naturally vanish. Thus, they are neglected when defining $a_{h}^{E}$ in the stabilized VC-VEM setting.
The theoretical analysis of VC-VEM involves several technical aspects and is beyond the scope of the present work.
Concerning the self-stabilized formulations $\mathrm{V1}$, $\mathrm{V3}$, $\mathrm{V4}$ and $\mathrm{V6}$, an higher-order counterpart of $\Pi^{\mathcal{A}}_{p}$ is computed within the framework of the discrete spaces defined in Section 3.2. Moreover, in lieu of the $L^{2}$ projection, the following operator is considered:
@@
the $\mathrm{VC}$-VEM does not apply to formulations $\mathrm{V2}$ and $\mathrm{V5}$ as the internal degrees of freedom would have the form $(\psi_{i}\,\text{div}\left(\mathcal{A}\left(x,y\right)\nabla q\right))_{E}$, which involves the possibly non-polynomial coefficient $\mathcal{A}\left(x,y\right)$.
@@ 7.2 Linear Elasticity problem
The VC-VEM is now extended to the linear elasticity problem defined in Eq. (42). As presented in the previous example, the elasticity space in Eq. (44) is considered. The following polynomial projection $\bm{\Pi}^{\mathbb{C}}_{p}\bm{\psi}_{i}$ of $\bm{\psi}_{i}$ ($i=1,...,N_{d}\big|_{E}$) is then employed to define the polynomial contribution to the discrete bilinear form:
$\begin{cases}\begin{aligned} &a^{E}\left(\bm{\Pi}^{\mathbb{C}}_{p}\bm{\psi}_{i},\bm{q}\right)=\mathfrak{a}_{\mathsf{vc}}^{E}\left(\bm{\psi}_{i},\bm{q}\right)&&\forall\bm{q}\in\left[\mathcal{P}_{p}\left(E\right)\right]^{2},\\ &P_{0}^{E}\left(\bm{\Pi}^{\mathbb{C}}_{p}\bm{\psi}_{i},\bm{q}\right)=P_{0}^{E}\left(\bm{\psi}_{i},\bm{q}\right)&&\forall\bm{q}\in RM\left(E\right),\end{aligned}\end{cases}$ (72)
where now:
$a^{E}\left(\bm{\psi}_{i},\bm{q}\right)=\int_{E}\bm{\varepsilon}\left(\bm{\psi}_{i}\right)^{T}\mathbb{C}\left(x,y\right)\bm{\varepsilon}\left(\bm{q}\right)\,\mathrm{d}E,$ (73)
and:
$\mathfrak{a}_{\mathsf{vc}}^{E}\left(\bm{\psi}_{i},\bm{q}\right)\coloneqq$ $-\int_{E}\bm{\Pi}^{\bm{0}}_{p}\bm{\psi}_{i}\,\bm{L}\left(\mathbb{C}\left(x,y\right)\bm{\varepsilon}\left(\bm{q}\right)\right)\;\,\mathrm{d}E$ (74)
$+\sum_{e=1}^{n_{e}}\int_{\Gamma^{E}_{e}}\bm{\psi}_{i}\cdot\hat{\bm{\sigma}}\left(\bm{q}\right)\bm{\hat{n}}_{\Gamma_{e}}\;\,\mathrm{d}\Gamma^{E}_{e}$
$+\sum_{e=1}^{\tilde{n}_{e}}\int_{\tilde{\Gamma}^{E}_{e}}\bm{\psi}_{i}\cdot\hat{\bm{\sigma}}\left(\bm{q}\right)\bm{\hat{n}}_{\tilde{\Gamma}_{e}}\;\,\mathrm{d}\tilde{\Gamma}^{E}_{e}.$
The symbol $\hat{\bm{\sigma}}$ denotes the Cauchy stress tensor accounting for the spatial variability of $\mathbb{C}\left(x,y\right)$.
@@
The projection operators $\Pi^{\mathcal{A}}_{p}$ and $\bm{\Pi}^{\mathbb{C}}_{p}$ must be recomputed whenever the coefficient changes. This also applies when a mesh of uniform elements is considered, as the coefficient varies spatially. In principle, the $\mathrm{VC}$-VEM may result in a higher computational cost. However, the overhead is mitigated by the fact that the $p$-version of VEM is generally defined on coarse meshes.
@@ 8.1.1 Example 1: polynomial coefficients
First, the quadrilateral mesh is considered. The results are shown in Figure 15. The accuracy of the solution, as seen from Figure 15(a), is the same for all the simulations up to the order $p=4$. However, when $p>4$ some curves exhibit a diverging response. It is worth noting the excellent accuracy displayed by all the methods based on the newly introduced approach to consider coefficient-variability, even for $p\geq 5$. In this scenario, also the self-stabilized $\mathrm{V3}$ and the stabilized VEM with $\Pi^{0}_{p-1}$ projection demonstrate excellent convergence. This behavior is attributed to the use of the $L^{2}$ projection, which tends to exhibits better accuracy in the presence of variable coefficients, as already discussed in [7]. On the other hand, the formulation $\mathrm{V6}$, which does not consider variable coefficients, behaves as the stabilized VEM with $\Pi^{\nabla}_{p}$. Regarding the stiffness matrix conditioning, reported in Figure 15(b), no differences can be noted between the cases with or without variable coefficient projection.